. Dry air will support a maximum electric field strength of about 2.0 x 108 V/m. Above that value, the field creates enough ionization in the air to make the air a conductor. This allows a discharge or spark that reduces the field. What, then, is the maximum voltage between two parallel conducting plates separated by 3.5 cm of dry air?


1
Expert's answer
2022-03-20T18:50:46-0400

The maximum voltage between two parallel conducting plates 

"V=Ed\\
=\rm 2.0\times 10^8\: V/m\times 0.035\: m=7.0\times 10^6\: V"